

********************************************************************
UARS WEEKLY STATUS REPORT - 31 Dec 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HAPPY NEW YEAR!!!!
The new PEM2K:: is now fully configured and functional. All communication
paths should function as normal.
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004199 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004200 and 20041229 has now been processed and waiting for archival
at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004199. MEPS data between 2004200 and 2004229
is currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 December 2004 (Orbit 72624, GMT Day 353) through
24 December 2004 (Orbit 72728, GMT Day 359).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 353/2314 72638 ZEPS/NEPS DAY-ONLY
PEM 356/2124 72682 ZEPS/NEPS FULLTIME
WINDII 357/2301 72698 POWER ON
WINDII 358/0142 72700 POWER OFF
The beta angle ranged between the angles of 63.4 degrees to 78.3
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 2.0 degrees on 12 January 2005 (DOY 012).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.94 degrees C. Additional
battery performance data for each battery on 03 December and 10
December 2004 are provided in Appendix D.
The clock error ranged from 10.8 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
PEM = 02
WINDII = 02
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week UARS flew through its winter high beta period. During this
time the spacecraft night is reduced to 45 minutes, which in turn
reduces the depth of discharge on the batteries. This also causes the
charge to discharge ratios to increase. In order to keep this ratio
down the FOT turned PEM ZEPS/NEPS on two days early and temporarily
turned on the WINDII pulse and quiet busses. The FOT will continue to
monitor the battery telemetry closely UARS comes back out of the high
beta range.
The FOT continued their CMS software acceptance testing this week.
This testing will continue into the new year. The goal for completion
is no later than end of January 2005.
The UARS FOT would like to extend their best wishes for this holiday
season to all those who helped make this year another successful year
for UARS.
MEETINGS
The FOT attended the biweekly meeting with the HRDI team to discuss
the attitude error seen in their data. The FDF team is closing in on
a final solution and the simulations on the UTTS continue to be run.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 January 2005 (DOY 012).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 December 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72633 353/1357 04-350 SUSIM EXCEPTION CLEAR
72638 353/2314 04-351 PEM ZEPS/NEPS DAY-ONLY
72663 355/1428 04-352 MPS VT L-6 TO L-5
72663 355/1428 04-353 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
72667 355/2017 04-354 MPS BATSOC COMMAND
72682 356/2124 04-355 PEM ZEPS/NEPS FULLTIME
72693 357/1420 04-356 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
72698 357/2301 04-357 WINDII POWER ON (POWER MNGMT)
72700 358/0142 04-358 WINDII POWER OFF
72723 359/1451 04-359 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 December 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 December 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72663 355/1428 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
72693 357/1420 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
72723 359/1451 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 December 2004
17 December 2004 - GMT Day 352 Beta = 61.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.99/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/ -56.0 +4.50/+3.23 +21.6/-13.6 26.9 1.05 17.0
3 -33.6/-140.0 +2.91/+1.97 +20.0/-12.8 26.7 1.04 16.3
24 December 2004 - GMT Day 359 Beta = 79.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.48/-2.42 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/ -44.8 +2.28/+1.65 +11.6/-12.8 27.7 1.08 9.6
3 -72.8/-184.8 +1.65/+0.71 +11.2/-12.0 27.7 1.07 9.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
72663 355/1428 VT L-6 TO L-5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 December 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-092 FOT 72635 353/1831 37983 Late Acq. - TDRS West
04-093 FOT 72717 359/0407 37991 No Acq. - TDRS 275
04-090 FOT 72719 359/0748 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-090 FOT See List Dec 04 30832 Generic Late Acquisition
04-091 FOT 72530 346/1854 37847 No UPD/GCMR Capability
04-092 FOT 72635 353/1831 37983 Late Acq. - TDRS West
04-093 FOT 72717 359/0407 37991 No Acq. - TDRS 275